OpenStudy (anonymous):

Problems for fun: How many numbers from 1- 200 aren't divisible by 2,3,5, or 7? I'll post a solution when my movie is over.

jhonyy9 (jhonyy9):

- the prime numbers

myininaya (myininaya):

i think i got it

myininaya (myininaya):

please hold ...

myininaya (myininaya):

so there are 180 numbers divisible by some combination of the above so 20 aren't

OpenStudy (anonymous):

i guess i should figure out the answer now >.< one sec lol

OpenStudy (anonymous):

Half of the numbers are even, hence divisible by 2. Of the remaining numbers, one third are divisible by 3, so we have 133 numbers divisible by 2 or 3. 40 numbers are divisible by 5, of these half are even so we have 20 numbers divisible by 5 and not by 2. 13 numbers of the original 200 are divisible by 3 and 5 (by 15) and of these 7 are odd and 6 even (we know there's one more odd number since 15 is odd) so 7 numbers are divisible by 3 and 5, but not by 2, so of the 20 numbers we had 6 are divisible by 3 and we find that there are 20 - 6 = 14 numbers that are divisible by 5 and not by 3 nor 2. Combining this with the previous result, there are 133 + 13 = 146 numbers that are divisible by 2, 3 or 5. 28 numbers are divisible by 7, half of which are even so we have 14 numbers divisible by 7 and not by 2. 5 numbers of the original 200 are divisible by 5 and 7 (by 45) and of these 3 are odd and 2 are even, so we have 14 - 3 = 11 numbers that are divisible by 7 and not by 2 nor 5. 9 numbers are divisible by 3 and 7 (by 21) 5 of which are odd and one of which is also divisible by 5 (by 105), 11 - 5 + 1 = 7 numbers are divisible by 7 and not by 2, 3 nor 5. Finally, there are 146 + 7 = 153 numbers that are divisible by 2, 3, 5 or 7 between 1 and 200, so there are exactly 47 numbers which are not.

myininaya (myininaya):

oops i need to go further with my approach

myininaya (myininaya):

i got 33 i think

myininaya (myininaya):

the rest

OpenStudy (anonymous):

im getting 47

OpenStudy (anonymous):

i didnt realize how long this would take to calculate <.< i thought 3 numbers was too short lol.

OpenStudy (anonymous):

scanning now~ btw nice job Krebante :)

OpenStudy (anonymous):

i caught a mistake in mine thanks you myininaya's pic, i had missed 2, 5, 7 lol

OpenStudy (anonymous):

thank's to*

myininaya (myininaya):

if you cant open my file there are 100 multiplies of 2 there are 66 multiplies of 3 there are 40 multiplies of 5 there are 28 multiplies of 7 there are 33 multiplies of 2*3 there are 20 multiplies of 2*5 there are 14 multiplies of 2*7 there are 13 multiplies of 3*5 there are 9 multiplies of 3*7 there are 5 multiplies of 5*7 there are 6 multiplies of 2*3*5 there are 4 multiplies of 2*3*7 there are 2 multiplies 2*5*7 there is 1 multiple of 3*5*7 so we 234 from adding the n(multiplies of 2)+n(multiplies of 3)+n(multiplies 5)+n(multiplies of 7) so we have some overlapping so need to subtract any intersections 234-n(multiplies of 2*3)-n(multiplies of 2*5)-n(multiplies of 2*7) - n(multiplies of 3*5)-n(multiplies of 3*7)-n(multiplies of 5*7)-n(multiplies of 2*3*5)-n(multiplies of 2*3*7)-n(multiplies of 2*5*7)-n(multiplies of 3*5*7) giving us 167 200-167=33 that are not divisible by 2,3,5,or 7

myininaya (myininaya):

what did i miss?

myininaya (myininaya):

ok i will look at yours lol i didn't realize u posted that

OpenStudy (anonymous):

in your method, you have 234 from adding the multiples of the "singles", then you subtract the multiples of the "doubles", now you need to add the multiples of the "triples". so 234-(33+20+14+13+9+5) = 140 Then you add: 140+(6+4+2+1) = 153 So its 200 - 153 = 47

myininaya (myininaya):

why do we add the triples?

myininaya (myininaya):

oh wait just in case we took out to much from the doubles i think

myininaya (myininaya):

my scan was ugly lol

OpenStudy (anonymous):

Even beyond the over counting we are doing in counting the singles, we are also over counting the doubles. for example, 140 is both a multiple of 5*2 and 2*7. So we need the triples to be the yeah you got it :)

myininaya (myininaya):

yours was alittle bit more pretty its hard for me to think unless i write down the way i think

OpenStudy (anonymous):

you just keep switching the sign from positive to negative (or vice-versa) to take care of over counting. So if i had said, divisible by 2,3,5,7,and11, then it would be TOTAL - (singles)+(doubles)-(triples)+(quadruples)-(quintuples)

myininaya (myininaya):

right i get it

OpenStudy (anonymous):

i only write nice when i know people are going to look at it lol, otherwise its chicken scratch >.< i can read it while im writing it, but say a day later, its like hieroglyphs lol

myininaya (myininaya):

nice problem i liked it

OpenStudy (anonymous):

its interesting! infact, this can be used to tell you if there are any more prime numbers in that range. since there were 47 numbers left out of the 200, that means there are at most 47 primes left. once i find another one i can reduce that number further. so on and so forth

myininaya (myininaya):

at first i was going to act like i didn't see your problem but then i was like oh maybe i should look at the opposite of what you said

OpenStudy (anonymous):

lol

myininaya (myininaya):

i was thinking about it because i didn't want to let you down though

myininaya (myininaya):

im going to try to get my hours corrected i gotta go to bed peace

OpenStudy (anonymous):

lm glad you tried it out :) regardless of if one gets it right or wrong, it encourages thought.

OpenStudy (anonymous):

me too, although im gonna get food lol >.< good night :)

myininaya (myininaya):

well i was on the right path at least and i had all the numbers there i just didn't do the operations right

myininaya (myininaya):

and at first i did forget some of my intersections

myininaya (myininaya):

namely the triples

myininaya (myininaya):

goodnight

OpenStudy (anonymous):

night :)